OpenStudy (anonymous):

just wanna make sure one more time :) Use the definition of f′(a) (derivative) to compute f′(a) when f(x) =4/(x+1) and a = 2: * * which one should i use : limX-->a = f(x)-f(a)/X-a or lim h-->0 = f(a+h)-f(a)/h

OpenStudy (unklerhaukus):

\[f(x)=\frac 4{x+1}=4(x+1)^{-1}\] \[f'(x)=-4(x+1)^{-2}\]